OpenStudy (anonymous):

You are a student who plans to attend the amusement park for two days. This will be fun! Now it's time to figure out the cost. You won a coupon for 50% off a 1-day pass but it cannot be combined with other offers. The amusement park also offers a $5 student discount; however, you cannot get the $5 student discount off the 2-day pass. Use the information in the chart above to determine which option is the better deal. • Option 1: Buying a ticket at 50% off today and then purchasing a new 1-day ticket at the student price tomorrow. • Option 2: Buying a 2-day pass and not using the coupon at all.

OpenStudy (anonymous):

someone please help i feel so lost!!!

OpenStudy (anonymous):

is there a missing chart?

OpenStudy (anonymous):

1 day at the park for $44.95 2 days for $71.95 3 days for $94.95 Groups of 10 pay $400.00 for 1-day tickets. A family with 2 adults and 3 children under 12 will pay only $174.45 for 1-day tickets. Student discounts are $5.00 off for 1-day tickets. Coupons of 50% off admission cannot be combined with any other offer.

OpenStudy (anonymous):

thts tha chart sorry

OpenStudy (anonymous):

Option 1: .5(44.95) + (44.95-5) = $62.425 (half price on the first day + $5 off on the second day) Option 2: $71.95 (two day pass from the table) obviously option 1 is the better deal

OpenStudy (anonymous):

more questions based on this scenario?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A business wants to give each of its employees a free ticket to the amusement park and has budgeted $1200 for tickets. 1. Write and solve an inequality to find the maximum number of 1-day, adult tickets that can be bought. When you round your answer, remember that there is no such thing as "part" of a ticket. how do i do tht?

OpenStudy (anonymous):

i kno i saw i apologize i didnt see it thts y i deleted tha comment

OpenStudy (anonymous):

ok no prob

OpenStudy (anonymous):

idk how to do inequalities :/ so how wud i go about doing the thing i just sent

OpenStudy (anonymous):

if n is the number of 1-day tickets purchased, then solving the inequality\[44.95n \le1200\] for n (rounding down to the next lower integer) will give the max number of tickets that can be purchased without exceeding $1200. solving you would get\[n \le \frac{1200}{44.95} \approx 26.69\]therefor the max number of tickets that they can purchase is 26 tickets. Note: 26 tickets times $44.95 = $1168.70 and if they try to buy one more they would exceed the $1200 spending cap: 1168.70 + 44.95 = $1213.65, which confirms the 26 tickets max

OpenStudy (anonymous):

O.o woah....

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

ummm...honestly..?

OpenStudy (anonymous):

ummm...honestly..?

OpenStudy (anonymous):

no, please lie

OpenStudy (anonymous):

haha im just playn yea it makes sense

OpenStudy (anonymous):

me too :)

OpenStudy (anonymous):

mo' questions?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

nope ure awesome for helping me

OpenStudy (anonymous):

i was happy to do it... later

OpenStudy (anonymous):

peace